Text Question 3.4 Consider the demand function for processed pork in Canada Qd = 526.00-28p + 20pb + 3pc + 0.002Y The supply function for processed pork in Canada is Qs = 410.00 + 36p-60ph p is the price of pork Q is the quantity of pork demanded pb is the price of beef = $4 per kg Pc is the price of chicken -$3 per kg (measured in millions of kg peryear)Y is the income of consumers $12,500 Ph is the price of a hog $1.50 per kg Solve for the equilibrium price and quantity for pork. T he equilibrium price of pork is Sand the e quilibrium quantity of pork is million kg per year. Enter numeric responses using real numbers rounded up to two decimal places.)
Answer #1

plugging the given values of price in both equations , we have

Qd= 526-28P+20*4 +3*3 +0.002*12500 = 640- 28P

Qs= 410 + 36P-60*1.5 = 320 + 36P

in equilibrium Qd=Qs; so 640-28P=320 + 36P ; P= 5

and Q= 640-28*5 =500